### v1

@@ 4 Theoretical analysis
Theorem 3 provides the error bound for the centralized estimator $\sum_{l}\overline{M}_{l}/L$. The RHS of (3) indicates that the privacy parameters $q$ and $q^{\prime}$ affect the accuracy of the estimated eigenvectors. In addition, the influence of $q$ and $q^{\prime}$ on the final bound is not equivalent.
First, we observe that the privacy parameters $q,q^{\prime}$ appear in the numerators of the first two terms of (3) in the form $q\rho+1-q^{\prime}$. Considering that $\rho=o(1)$, we can see that a large level of $q^{\prime}$ is more desirable than a large level of $q$ for reducing the estimated error (3). In other words, perturbing 0 to 1 is more harmful than perturbing 1 to 0 for the model estimation. Second, when $\rho\leq C\frac{(q\rho+1-q^{\prime})^{1/2}}{(q+q^{\prime}-1)}\cdot\frac{\log^{1/2}(L+n)}{n^{1/2}}$, the RHS of (3) reduces to
@@
$\max\left\{\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\right\},$
which matches the following bound without privacy-preserving up to log factors derived in Lei and Lin (2022),
$\max\left\{\frac{\rho\log^{1/2}(L+n)}{\sqrt{L}},\,\rho^{2}\right\},$ (15)
provided that $q$ and $q^{\prime}$ both approach to 1.
The derivation of the bound for the single matrix $\overline{M}_{l}$ is similar to that in Theorem 3. For the sake of space, we summarize the results in Theorem B.1 in the Appendix.
Now we are ready to specify the unknown terms in Theorem 2.
@@
Suppose Assumptions 1-4 and condition (13) hold. If
$\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(n)}{n}}\quad{and}\quad\mathcal{H}\lesssim n\rho^{2},$ (16)
then the output $\tilde{V}$ of Algorithm 1 satisfies
$\mathrm{dist}(\tilde{V},V)\lesssim\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}$ (17)
with
$\mathcal{I}_{1}:=\frac{q\rho+1-q^{\prime}}{(q+q^{\prime}-1)^{2}}\cdot\frac{\log n}{n\rho^{2}},\quad\mathcal{I}_{2}:=\frac{\mathcal{H}^{2}(n,\rho,L)}{n^{2}\rho^{4}},$ (18)
and
$\small\mathcal{I}_{3}:=\max$$\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)n^{1/2}\rho\sqrt{L}},\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}n\rho^{2}\sqrt{L}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}n\rho^{2}}\bigg{\}},$ (19)
with probability larger than $1-O((L+n)^{-\nu})-O(Ln^{-\nu})$ for some constant $\nu>0$.
To make the upper bound (17) valid, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ both should be small enough. To this end, the technical condition (16) is utilized to meet the requirement in Theorem 2. The results in Corollary 1 can be simplified in various ways. For example, if we further require $\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(L+n)}{n}}$, then the first term in $\mathcal{I}_{3}$ dominate the second one. And the bound would be further reduced to
$\frac{(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)n^{1/2}\rho\sqrt{L}}$ (20)
if $\mathcal{I}_{2}$ is suppressed by $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$, and $L\lesssim\frac{(q+q^{\prime}-1)^{2}}{q\rho+1-q^{\prime}}\cdot\frac{n\rho^{2}\log(L+n)}{\log^{2}n}$. As expected, we see in (20) that more local networks (large $L$) and less stringent privacy-preserving demand (large $q$ and $q^{\prime}$) lead to a smaller projection distance.
With Corollary 1 at hand, it is convenient to derive the following misclassification error bound of ppDSC.
@@
Suppose the conditions in Corollary 1 all hold. Denote the set of misclassified nodes by $\mathcal{M}$. Then, with probability larger $1-O((L+n)^{-\nu})-O(Ln^{-\nu})$ for some constant $\nu>0$, Algorithm 1 has misclassification rate upper bounded as
$\frac{|\mathcal{M}|}{n}\leq(\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3})^{2}$
with $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ defined in (18) and (19).
Note that the misclassification error is just the squared error of the eigen-vectors $\tilde{V}$ in Corollary 1. Therefore, Theorem 4 can be similarly discussed and simplified. We omit the discussion for the sake of space.
Finally, we discuss the necessity of the bias-adjustment procedure in Appendix B. We investigate the deviation of the non-debiased matrix $\tilde{M}:=\frac{1}{L}\sum_{l}\frac{1}{n}\tilde{A}_{l}^{2}$ from population matrices $Q$ and $\tilde{Q}:=\frac{1}{L}\sum_{l}\left((q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}\right)^{2}/{n}$, respectively. It turns out that the bias-adjustment procedure can lead to a smaller error of eigenvector and misclassification, no matter which population matrix is used.
@@ 5 Simulations
In this section, we evaluate the finite sample performance of the proposed algorithm. We show how the privacy parameters, the number of networks, and the number of nodes affect the finite sample behavior of the resulting estimates. In Appendix C, we test the effect of the heterogeneity on the performance of the proposed algorithm.
@@
Our algorithm is denoted by ppDSC. We compare our algorithm, ppDSC, with the following algorithms.
  • ppDSC-1b: computed according to Algorithm 1 but only with one-step bias-correction (i.e., diagonal-debiasing) given in Lei and Lin (2022). Note that our bias-correction for the diagonal elements is not identical to Lei and Lin (2022);
  • ppDSC-2b: computed according to Algorithm 1 but without any bias-correction, which corresponds to Charisopoulos et al. (2021);
  • Oracle: a centralized algorithm. It is obtained from the spectral clustering based on the debiased sum of squared RR-perturbed matrices (i.e., $\overline{M}$ defined in (10)). Though violating the local computation constraint, this is used as a baseline algorithm.
@@
$\scriptsize B^{(1)}=W\begin{bmatrix}1.5&&\\ &0.2&\\ &&0.4\end{bmatrix}W^{\intercal}\approx\begin{bmatrix}0.62&0.22&0.46\\ 0.22&0.62&0.46\\ 0.46&0.46&0.85\end{bmatrix},$
and
@@
We evaluate the effect of $q,q^{\prime}$, $n$ and $L$ on the numerical performance. The parameter set-up is as follows.
  • Effect of $L$: $n=210$, $q=q^{\prime}=0.9$, and $L$ ranges from 2 to 20.
  • Effect of $n$: $L=10$, $q=q^{\prime}=0.9$, and $n$ ranges from 15 to 300.
  • Effect of $q,q^{\prime}$: $L=10$, $n=210$, and $q$ and $q^{\prime}$ are equal and range from 0.6 to 1.
Results.
The averaged results over 20 replications are displayed in Figure 1. We have the following observations. First, as expected, as $q,q^{\prime}$, $n$, and $L$ increase, the performance of all estimators improve. Second, our estimator ppDSC is superior to the biased ppDSC-1b and ppDSC-2b except when the privacy parameters $q,q^{\prime}$ become close to 1 and the sample size $n$ is large. It is because that ppDSC-1b and ppDSC-2b ignore the bias caused by privacy-preserving the networks and thus mis-specify the model. When $q,q^{\prime}$ become close to 1, the bias from privacy preservation does not matter much, and then ppDSC-1b performs similarly to ppDSC. Third, as with extra local computation demand, our method ppDSC is generally not better than the Oracle estimator. However, when $q,q^{\prime}$, $n$, and $L$ become large, they are getting close. This is consistent with our theoretical results.
@@ Results.
(f) Effect of $L$
Figure 1: Comparison of ppDSC, ppDSC-1b, ppDSC-2b and Oracle on the simulated data. The effect of the privacy parameters $q,q^{\prime}$, the number of nodes $n$, and the number of networks $L$ are shown, respectively. The projection distance and misclassification rate are evaluated.

### v2

@@ 4 Theoretical analysis
Remark 5
Condition (13) implicitly assumes that $q,q^{\prime}$ relate to the privacy budget $\epsilon$ via (4). Nonetheless, the proof of Theorem 3 indicates that the result (3) also holds under (3), i.e., the more general region of $q,q^{\prime}$, with a modification of condition (13). Theorem 3 provides the error bound for the centralized estimator $\sum_{l}\overline{M}_{l}/L$. When the privacy parameter $q^{\prime}$ is large in the sense that $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$, the upper bound in (3) matches that in Lei and Lin (2022) up to log terms, where they did not impose the privacy-preserving constraint.
Similar to Theorem 3, we can easily derive the upper bound for $\max_{i}\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/s\|_{2}$. Putting together the results, we provide in the following Theorem 4 the specific error bound for the estimated eigen-spaces by Algorithm 1.
It is well-known that large network density will help to recover the communities. Therefore, in what follows, we focus on the more challenging regime with $n{\rho^{2}}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\lesssim{{\log(s+n)}}$, though our results can also be extended to $n{\rho^{2}}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\gtrsim{{\log(s+n)}}$.
Theorem 4
Suppose Assumptions 1-3 and condition (13) hold. If $n{\rho^{2}}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\lesssim{{\log(s+n)}}$,
$n\sqrt{s}\cdot\frac{\rho^{2}}{C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}}}\gtrsim{{\log(s+n)}};$ (15)
$(q+q^{\prime}-1)\cdot\frac{\xi_{q,q^{\prime}}}{\rho}\lesssim n^{1/2},$ (16)
then the output $\tilde{V}$ of Algorithm 1 satisfies
$\mathrm{dist}(\tilde{V},V)\lesssim\mathcal{E}_{1}+\mathcal{E}_{2}+\mathcal{E}_{3};$
$\mathcal{E}_{1}:=\max\bigg{\{}\frac{\log^{2}(s+n)(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})^{2}}{n^{2}{s}\rho^{4}},\;\frac{(1+(q+q^{\prime}-1)\xi_{q,q^{\prime}}/\rho)^{4}}{n^{2}}\bigg{\}},\;\mathcal{E}_{2}:=\frac{\mathcal{H}^{2}(n,\rho,L)}{n^{2}\rho^{4}};$
$\mathcal{E}_{3}:=\max\bigg{\{}\frac{\log(L+n)(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})}{n{\sqrt{L}}\rho^{2}},\;\frac{(1+(q+q^{\prime}-1)\xi_{q,q^{\prime}}/\rho)^{2}}{n}\bigg{\}}$
with probability larger than $1-O((L+n)^{-\nu})-O(m(s+n)^{-\nu}))$ for some constant $\nu>0$.
In Theorem 4, $\mathcal{E}_{1},\mathcal{E}_{2}$ and $\mathcal{E}_{3}$ correspond to the squared local error, effective heterogeneity, and centralized error, respectively. The first two errors are the price to pay for the unavailability of full network layers. Condition (15) and (16) are required to make the results in Theorem 2 valid, which ensures $\mathcal{I}_{1}\lesssim 1$ in Theorem 2. $\mathcal{I}_{2}\lesssim 1$ is automatically met as we have shown in the proof. Several remarks are in order to explain the conditions and results in Theorem 4.
Remark 6 (Effect of network sparsity)
Condition (15) on $\rho$ depends on the number of network layers $s$ in each local machine. It reduces to $L^{1/2}n\rho\gtrsim\log(L+n)$ when $m=1$ and $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$, which is identical to the condition in Theorem 1 of Lei and Lin (2022) up to log terms. It is worth mentioning that the squared local error $\mathcal{E}_{1}$ and the follow-up condition on $\rho$ is inherent in the distributed learning algorithm with only one round of communication. See Zhang et al. (2013); Lee et al. (2017) for example in the context of distributed empirical risk minimization and distributed sparse regression. In the context of SBMs, as far as we are aware, this work is the first result to obtain a regression-like result for the one-shot-based distributed algorithm. It is also worth mentioning that Arroyo et al. (2021) studied community detection on multi-layer networks, with each layer distributively stored in a separate machine. The requirement on $\rho$ turns out to be $\Omega(\log n/n)$, independent of network layers, where $n$ denotes the number of nodes. The condition on $\rho$ may be alleviated if we consider multi-round communications as in the context of regression; see Jordan et al. (2018); Chen et al. (2020), among others. The power method-based approach provides a possible basis Guo et al. (2024), which we left as our future work.
Condition (15) and (16) both reveal that denser networks (i.e., larger $\rho$) can allow a more strict privacy-preserving level (i.e., small $\epsilon$). To see this, suppose $q,q^{\prime}$ relate to $\epsilon$ via (4), then (15) reduces to $\gamma(\epsilon):=\frac{(\mathrm{e}^{\epsilon}-1)^{2}}{\mathrm{e}^{\epsilon}+1}\gtrsim\frac{1}{\rho^{2}}\cdot\frac{\log(s+n)}{{ns^{1/2}}}$, where $\gamma(\epsilon)$ increases with $\epsilon$; (16) reduces to $\mathrm{e}^{\epsilon}-1\gtrsim{1}/{(\rho n^{1/2})}$.
Remark 7 (Asymmetric roles of $(q,q^{\prime})$)
From the definition of $C_{q,q^{\prime}}$ and $\xi_{q,q^{\prime}}$, it is undoubtedly that in general, the error bound of the eigenspace decreases when $q$ and $q^{\prime}$ values increase. Interestingly, we find that without a pre-specified privacy budget $\epsilon$, $q$ and $q^{\prime}$ play an asymmetric role in the error bound when the network is sparse with $\rho=o(1)$. The asymmetric role can be observed through the term $C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}}$. When $q,q^{\prime}>1/2+\tau$ for an arbitrarily small constant $\tau>0$, $C_{q,q^{\prime}}$ is a constant and $\xi_{q,q^{\prime}}\asymp 1-q^{\prime}$. Thus, $C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}}\asymp\rho+1-q^{\prime}$, so we see that larger $1-q^{\prime}$ would make the error bound larger, i.e., perturbing 0 to 1 is more harmful to the upper bound than perturbing 1 to 0. This interesting phenomenon can be explained as follows. We see that the sparsity level of the RR-perturbed network has the order of $\rho+1-q^{\prime}$, so perturbing 1 to 0 does not change the order of the network sparsity so it does not affect the error bound with a large sample. However, perturbing 0 to 1 can make the network denser if $1-q^{\prime}>\rho$, but the inflated “1” values are the wrong and harmful ones, making the error bound worse.
Remark 8 (Optimal choice of $(q,q^{\prime})$)
With a pre-specified specified privacy budget $\epsilon$, Theorem 4 indicates that $q=q^{\prime}=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ is the best choice to yield the smallest statistical error bound for the eigenspace estimation. In particular, we can easily show the three terms $C_{q,q^{\prime}}$, $\xi_{q,q^{\prime}}$ and $(q+q^{\prime}-1)\xi_{q,q^{\prime}}$ involving $q,q^{\prime}$ all attain their minimum at $q=q^{\prime}=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$. For example, consider $\xi_{q,q^{\prime}}$, plugging in the upper boundaries shown in Figure 1, i.e., $q^{\prime}=1-\frac{q}{\mathrm{e}^{\epsilon}}$ when $q\leq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ and $q^{\prime}=\mathrm{e}^{\epsilon}(1-q)$ when $q\geq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$, we have
@@
$\xi_{q,q^{\prime}}=\frac{1-q^{\prime}}{(q+q^{\prime}-1)^{2}}=\left\{\begin{array}[]{ll}\frac{1}{q\mathrm{e}^{\epsilon}(1-\mathrm{e}^{-\epsilon})^{2}}&\text{if }q\leq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}},\\ \frac{1-\mathrm{e}^{\epsilon}+\mathrm{e}^{\epsilon}q}{(1-q)^{2}(1-\mathrm{e}^{\epsilon})^{2}}&\text{if }q\geq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}},\end{array}\right.$
which attains minimum at the $q=\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ and at this time $q^{\prime}$ equals $q$. The terms $C_{q,q^{\prime}}$ and $(q+q^{\prime}-1)\xi_{q,q^{\prime}}$ can be derived similarly.
The next corollary shows that under certain extra conditions, the bound of our distributed estimator reduces to its centralized counterpart.
@@
Suppose the assumptions and conditions in Theorem 4 all hold. If we further have
$m\lesssim\frac{n\rho^{2}}{(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\log(s+n)}\cdot L^{1/2};$ (17)
$\mathcal{H}\lesssim\max\bigg{\{}\frac{n^{1/2}\rho^{2}(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})^{1/2}\log^{1/2}(L+n)}{L^{1/4}\rho},\,{n^{1/2}\rho}(\rho+(q+q^{\prime}-1)\xi_{q,q^{\prime}})\bigg{\}},$ (18)
then $\mathcal{E}_{3}$ dominates $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$; see their definitions in Theorem 4. The output of $\tilde{V}$ of Algorithm 1 satisfies
$\mathrm{dist}(\tilde{V},V)\lesssim\max\bigg{\{}\frac{\log(L+n)(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})}{n{\sqrt{L}}\rho^{2}},\;\frac{(1+(q+q^{\prime}-1)\xi_{q,q^{\prime}}/\rho)^{2}}{n}\bigg{\}}$
with probability larger than $1-O((L+n)^{-\nu})-O(m(s+n)^{-\nu}))$ for some constant $\nu>0$.
We make the following remarks on Corollary 1.
Remark 9
Condition (17) requires the number of machines should not be large given other parameters. In particular, when ${n}\rho^{2}/(C_{q,q^{\prime}}\rho+\xi_{q,q^{\prime}})\asymp{{\log(s+n)}}$, (17) reduces to $m\lesssim\sqrt{L}$, which is a crucial condition for the distributed estimator to match the error rate of its centralized counterpart in one-shot-based distributed learning literature, see Zhang et al. (2013); Lee et al. (2017), among others. The requirement for $m$ might be alleviated if one considers multi-round communications; see Jordan et al. (2018); Chen et al. (2020) in the context of regression.
Condition (18) is mild. For example, suppose $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$ and $\rho\asymp\frac{\log(s+n)m}{n\sqrt{L}}$ (i.e., the boundary in (17)), then (18) requires $\mathcal{H}\lesssim\frac{m^{2}\log^{1/2}(L+n)\log^{3/2}(s+n)}{nL}$. On the other hand, we naturally have $\mathcal{H}\lesssim n\rho^{2}\asymp\frac{m^{2}\log^{2}(s+n)}{nL}$ (see the proof of Theorem 4).
Remark 10
The error bound is comparable to Lei and Lin (2022) when $\xi_{q,q^{\prime}}\lesssim C_{q,q^{\prime}}\rho$. In particular, if $\rho\asymp\frac{\log(s+n)m}{n\sqrt{L}}$ and $\epsilon\asymp\log(\frac{n\sqrt{L}}{m})$ (suppose $q,q^{\prime}$ relates to $\epsilon$ via (4)), then $\mathcal{E}_{3}$ reduces to $\max\{\frac{\log\,(L+n)}{n\rho\sqrt{L}},\,\frac{1}{n}\}$, which is identical (up to log terms) to that in Lei and Lin (2022). Here, we remark that the large $\epsilon$ in the context of local-DP (Duchi et al., 2018) (i.e., DP at the level of original data rather than aggregated data) is reasonable and has also been used in other literature. We note that Mohamed et al. (2022) chose $\epsilon\gtrsim\log\,n$ when they studied the strong consistency of community detection algorithms under RR-perturbed SBMs. We also note that the work by Bhowmick et al. (2018) revisited the protections of local-DP (Duchi et al., 2018) and it is said that algorithms that only protect against the accurate reconstruction of (functions of) an individual’s data could allow a larger privacy budget ($\epsilon\gg 1$).
The following theorem shows that the misclassification error bound for ppDSC turns out to be the squared error of the eigen-vectors $\tilde{V}$.
@@
Suppose the conditions in Corollary 1 all hold. Denote the set of misclassified nodes by $\mathcal{M}$. Then, with probability larger $1-O((L+n)^{-\nu})-O(m(s+n)^{-\nu}))$ for some constant $\nu>0$, Algorithm 1 has misclassification rate upper bounded as $\frac{|\mathcal{M}|}{n}\lesssim\mathcal{E}_{3}^{2};$ see the definition of $\mathcal{E}_{3}$ in Theorem 4.
Finally, we discuss the necessity of the bias-adjustment procedure in Section B of the Supplementary Materials. We investigate the deviation of the non-debiased matrix $\tilde{M}:=\frac{1}{L}\sum_{l}\frac{1}{n}\tilde{A}_{l}^{2}$ from population matrices $Q$ and $\tilde{Q}:=\frac{1}{L}\sum_{l}\left((q+q^{\prime}-1)P_{l}+(1-q^{\prime})\textbf{1}_{n}\textbf{1}_{n}^{\intercal}\right)^{2}/{n}$, respectively. It turns out that the bias-adjustment procedure can lead to a smaller eigenvector and misclassification error bound, no matter which population matrix is used.
@@ 5 Simulations
In this section, we evaluate the finite sample performance of the proposed algorithm. We show how the number of networks, the number of nodes, and the number of local machines affect the finite sample behavior of the resulting estimates. We also justify our theory in terms of how to choose the edge-flipping probabilities in RR under a given privacy budget and how the network sparsity affects the performance under different privacy budgets. In Section C of the Supplementary Materials, we test the effect of heterogeneity and network sparsity on the performance of the proposed algorithm.
@@
Our algorithm is denoted by ppDSC. We compare our algorithm with the following two distributed algorithms denoted by ppDSC-1b and ppDSC-2b, both of which follow the above practical restrictions. Hence, ppDSC-1b and ppDSC-2b suffer from the model misspecification problem. We also compare two baseline centralized algorithms. One is the centralized counterpart of the proposed algorithm, denoted by ppSC. The other is the algorithm in Lei and Lin (2022), denoted by Oracle because it knows the unperturbed network like an oracle.
  • ppDSC-1b: computed according to Algorithm 1 but with line 6 replaced by one-step bias-correction (i.e., diagonal-debiasing) given in Lei and Lin (2022) (i.e., $\tilde{A}_{l}^{2}-\mbox{diag}(\tilde{A}_{l}^{2})$);
  • ppDSC-2b: computed according to Algorithm 1 but without any bias-correction, which corresponds to Charisopoulos et al. (2021) with RR-perturbed networks as input;
  • ppSC: obtained from the spectral clustering based on the debiased sum of squared RR-perturbed matrices (i.e., $\overline{M}$ defined in (9)).
  • Oracle: the algorithm in Lei and Lin (2022); obtained from the spectral clustering based on the diagonal-debiased sum of squared original adjacency matrices (i.e., $\sum_{l}(A_{l}^{2}-{\rm diag}(A_{l}^{2})))$.
@@
$\scriptsize B^{(1)}=W\begin{bmatrix}1.5&&\\ &0.2&\\ &&0.4\end{bmatrix}W^{\intercal}\approx\begin{bmatrix}0.62&0.22&0.46\\ 0.22&0.62&0.46\\ 0.46&0.46&0.85\end{bmatrix};$
@@
Experiment 1.
We evaluate the effect of $L$, $n$ and $m$ on the numerical performance. The parameter setup is as follows.
  • Effect of $L$: $n=210$, $q=q^{\prime}=0.8$, $m=L$, and $L$ from 2 to 17.
  • Effect of $n$: $L=m=12$, $q=q^{\prime}=0.8$, and $n$ ranges from 60 to 230.
  • Effect of $m$: $L=12$, $n=210$, $q=q^{\prime}=0.8$, and $m$ varies in $\{2,4,6,12\}$.
The averaged results over 20 replications are displayed in Figure 2. We have the following observations. First, as expected, as $L,n$ increases and $m$ decreases, the performance of all estimators improves. Second, our estimator ppDSC is superior to the biased ppDSC-1b and ppDSC-2b. It is because ppDSC-1b and ppDSC-2b ignore the bias caused by privacy-preserving the networks and thus mis-specify the model. Third, as with extra local computation and privacy preservation demand, our method ppDSC is generally not better than the centralized methods ppSC and Oracle. However, when $L$ and $n$ become large, and $m$ becomes small, they are getting close. This is consistent with our theoretical results.
@@
(f) Effect of $m$
Figure 2: Comparison of ppDSC, ppDSC-1b, ppDSC-2b, ppSC and Oracle on the simulated data. The effect of the number of networks $L$, the number of nodes $n$, and the number of local machines $m$ are shown, respectively. The projection distance and misclassification rate are evaluated.
Experiment 2.
We evaluate the performance of ppDSC with the privacy budget $\epsilon=1$ fixed and the edge-fillping probabilities $q,q^{\prime}$ varying in the feasible region of DP shown in Figure 1. Other parameters are $n=210$ and $L=m=12$. The averaged misclassification rates over 10 replications under different couples of $q$ and $q^{\prime}$ are displayed in Figure 3, where the infeasible region is filled by pure white. When $q,q^{\prime}$ are getting close to the upper boundaries, the performance of ppDSC becomes better. In particular, the misclassification rate attains the empirical minimum at the position marked by the pink cross, which is quite close to the theoretically best choice of $q=q^{\prime}=\frac{{\rm e}^{\epsilon}}{1+{\rm e}^{\epsilon}}$ (see the discussions after Theorem 3) marked by the red star.
Figure 3: The clustering performance of ppDSC with a pre-specified $\epsilon=1$ and $q,q^{\prime}$ varying in the feasible region of DP (see Figure 1). The empirically and theoretically best combination of $q,q^{\prime}$ are marked with red star and pink cross, respectively.